A bullet of mass m is fired from a gun of mass M .The recoiling gun compresses a spring of force constant K by a distance d. Then the velocity of the bullet in terms of K,d,M and m will be

Grade:12th pass

2 Answers

Khimraj
5 years ago
Let velocity of bullet is v and velocity of gun is V. ThenBy conservation of momentumMV = mvAnd as
Khimraj
5 years ago
MV = mv
and as given
½MV2 = ½Kd2
V = d\sqrt{K}
So v = (Md/m)\sqrt{K}
Hope it clears. If you like answer then please approve it.